# DomainForensics: Exposing Face Forgery across Domains via Bi-directional Adaptation

### v1

@@ IV-C Results of Cross-manipulation-datasets
[…]
TABLE V: Performance comparison of our method with state-of-the-arts under cross-faceswap&GAN scenario from FF++ to StyleGAN.
@@ IV-D Results of Cross-manipulation-types
In this part, we investigate the feasibility of our method to adapt faceswap faces to GAN-synthesized faces. Faceswap replaces the central region of face and retains other regions unchanged, while GAN synthesizes the whole face image. We use each of faceswap methods in FF++ [13] as source and adapt it to the StyleGAN dataset [58]. Table V shows the performance of our method in comparison to the three benchmark methods, MATT [61], RECCE [62] and SBI [18], under this setting. For fair comparison, we use the codes provided in [61, 62, 18] for SBI, MATT and RECCE under their defaulting settings. Observe that these three methods perform poorly, and in most cases their performance are below 50%. This is because these methods are designed to detect faceswap DeepFakes, e.g., finding the blending artifacts, and they cannot handle StyleGAN synthesized faces. In contrast, our method outperforms these methods by a large margin, on average 37%, 34% and 24% better at HQ level, and 117%, 20% and 10% better at LQ level, than MATT, RECCE and SBI, respectively. This is because our method focuses on learning the common forgery knowledge in the central regions of both faceswap and GAN DeepFakes, instead of simply exposing the faceswap-specific forgery features.
@@ IV-E1 Various Adaptation Settings
To provide more insights, we further investigate the proposed bi-directional adaptation strategy on the LQ set of FF++. We use the DeepFake detector trained on the source domain without adaptation as the baseline. GRL [70] and MMD [44] are two classical domain adaptation methods. GRL uses the gradient reversal layer and MMD attempts to reduce the distance of the probability distributions between source and target domains. GRL is exactly the domain adaptation method adopted in FEAT [54]. We denote the forward adaptation method in our method as FA. In the top part of Table VI, we evaluate these three domain adaptation methods. It can be seen that adding GRL, MMD or FA can improve the performance in the target domain, and using our FA attains the highest performance gain.
TABLE VI: Effect of various adaptation settings.
@@
Then we take a further step to analyze the effect of backward adaptation in the bottom part of Table VI, where SD means that the self-distillation is used in backward adaptation, while Ent. represents that self-distillation is replaced with entropy minimization of samples, which is inspired by the method of [71]. We observe that only adding SD can notably improve the performance on the target domain, but its performance on the source domain is compromised, especially in NT $\rightarrow$ FS, with $77.53\%$ of +SD v.s. $87.61\%$ of Baseline. This is because self-distillation is effective on the target domain, but overlooks the adaptation learned in forward adaptation. It can also be seen that replacing SD with Ent. is a bad idea, as this strategy degrades the detection performance in both domains. By adding both our forward adaptation and backward adaptation components to the baseline, our method can reach a good balance of the detection performance in both source and target domains. This clearly demonstrates that the proposed bi-directional adaptation is highly effective.
@@ IV-E2 Various Amounts of Training Samples
This part investigates the data efficiency of the proposed method. We randomly select a proportion or percentage of the samples in target domain for training to simulate data-constrained scenarios. Fig. 7 depicts the performance of our method for the adaptation pairs of DF $\rightarrow$ F2F, DF $\rightarrow$ NT, and F2F $\rightarrow$ FS on the LQ set of FF++ using the percentage of the target samples ranging from 20% to 100%. As expected, increasing the percentage generally improves achievable performance. More importantly, noting the baseline+FA performance of 86.49 for DF $\rightarrow$ F2F from Table VI, it can be seen that our method can still considerably enhance the performance to over 90 even if only 20% of target samples are used. This indicates that our method can be utilized in the more restricted cases where the available target samples are insufficient.
TABLE VII: Effect of frequency module.
@@ IV-E3 Effect of Frequency Module
The number of channels (C) and depth (D) of DCT blocks are two important factors for extracting frequency features. The C and D values used in our method are 768 and 4. We now investigate the achievable performance of our method using various C and D values on the LQ set of FF++. As can be seen from Table VII, given depth $\text{D}=4$, increasing channels from $\text{C}=192$, 384 to 768 improves the achievable performance on both source and target domains. Likewise, with channels fixed to $\text{C}=768$, increasing depth from $\text{D}=2$, 3 to 4 improves the achievable performance on both domains. The results of Table VII also validate that our choice of $\text{C}=768$ and $\text{D}=4$ is appropriate.
TABLE VIII: Performance of using various feature extractors.
@@ IV-E4 Various Feature Extractors
As mentioned in the introduction section, the proposed framework is independent of the architecture. Table VIII shows the performance of applying our framework on ResNet-50 [30], Xception [31], EfficientNet-B0 [32], EfficientNet-B4 [32], ViT-small [29] and ViT-base [29] on the LQ set of FF++. It can be seen that for each base network, our method significantly improves the performance on the target domain while maintaining a favorable performance on the source domain. This demonstrates that our proposed method is generically applicable.
@@ IV-E5 Added on State-of-the-art DeepFake Detection
This part investigates the effectiveness of our method to improve recently developed advanced detection methods. We use SBI [18] as an example. Specifically, we retrain SBI with HQ and LQ faces in FF++, and compare it with our method added on SBI. In this experiment, the source domain consists of fake faces blended by SBI and pristine faces and the target domain is each manipulation method. The results shown in Table IX reveal that with our method added on, SBI notably improves the performance both at HQ and LQ levels. This indicates that our method can be easily integrated into other detection methods to enhance their generalization ability.
TABLE IX: Performance of adding our method on SBI.
Method DF F2F NT FS Avg.
(HQ) SBI [18] 85.88 80.07 72.47 75.62 78.51
(HQ) SBI + Ours 99.00 97.23 90.28 92.92 94.86
(LQ) SBI [18] 80.70 67.63 65.13 62.50 68.99
(LQ) SBI + Ours 88.93 68.20 64.98 65.52 71.91

### v2

@@ IV-C Results of Cross-manipulation-datasets
In this experiment, we use FF++ as source domain and Celeb-DF as target domain. Table IV compares our method with the state-of-the-art methods, F3-Net [19], MATT [61], RECCE [62], Two-branch [66], LTW [34], DSP-FWA [15], MTD-Net [36], Xception [13], CFFs [37], DAM [67], SOLA-sup [63], SPSL [20], LiSiam [38], SLADD [35], LipForensics [68], HFI-Net [39], FTCN [64], AltFreezing [72], TALL [73], F2Trans [40], and PCL+I2G [17]. It can be seen from Table IV that our method achieves $90.22\%$ at HQ and $81.39\%$ at LQ, respectively, outperforming most other methods by a large margin. Note that PCL+I2G [17] synthesizes forged faces with raw quality data since the uncompressed data can provide more distinct features than the compression ones. Our method outperforms TALL-ViT [73], which uses spatio-temporal modeling for learning local and global contextual deepfake patterns, when employing a vision transformer as the backbone network. Additionally, it achieves comparable performance with TALL-Swin. This demonstrates that our method is capable of transferring the forgery knowledge from one dataset to another. The last two columns of Fig. 6 depict the examples of Grad-CAM from FF++ to Celeb-DF, which shows that our method concentrates more on forgery regions than the case without using our method.
For a comprehensive study, we then conduct a more challenging scenario, which uses the FF++ as the source domain and the DFDCP [71] or FFIW [67] as the target domain. The results are presented in Table V. We observe that our method achieves better performance than others on the DFDCP dataset, mainly because our method can learn common forgery features among different manipulated methods. However, SBI surpasses our method on the FFIW dataset. This is because this dataset contains many side faces, while the faces in FF++ dataset are usually frontal. This discrepancy largely increases the domain gap, which may disturb our model in learning common forgery features. Since SBI is designed to create a variety of faces for data augmentation, which likely includes such distorted faces in training, resulting in better performance than us.
TABLE V: Performance comparison of our method with state-of-the-arts under cross-datasets scenario from FF++ to DFDCP and FFIW.
Method Training Set DFDCP [71] FFIW [67]
PCL+I2G [17] FF++(RAW) 74.37 -
SRM [21] FF++(HQ) 79.70 -
SBI [18] FF++(RAW) 86.15 84.83
SBI [18] FF++(HQ) 85.51 83.22
Baseline FF++(HQ) 80.37 63.85
Baseline+BA (Ours) FF++(HQ) 89.57 76.76
TABLE VI: Performance comparison of our method with state-of-the-arts under cross-faceswap&GAN scenario from FF++ to StyleGAN.
@@ IV-D Results of Cross-manipulation-types
In this part, we investigate the feasibility of our method to adapt faceswap faces to GAN-synthesized faces. Faceswap replaces the central region of face and retains other regions unchanged, while GAN synthesizes the whole face image. We use each of faceswap methods in FF++ [13] as source and adapt it to the StyleGAN dataset [58]. Table VI shows the performance of our method in comparison to the three benchmark methods, MATT [61], RECCE [62] and SBI [18], under this setting. For fair comparison, we use the codes provided in [61, 62, 18] for SBI, MATT and RECCE under their defaulting settings. Observe that these three methods perform poorly, and in most cases their performance are below 50%. This is because these methods are designed to detect faceswap DeepFakes, e.g., finding the blending artifacts, and they cannot handle StyleGAN synthesized faces. In contrast, our method outperforms these methods by a large margin, on average 37%, 34% and 24% better at HQ level, and 117%, 20% and 10% better at LQ level, than MATT, RECCE and SBI, respectively. This is because our method focuses on learning the common forgery knowledge in the central regions of both faceswap and GAN DeepFakes, instead of simply exposing the faceswap-specific forgery features.
@@ IV-E1 Various Adaptation Settings
To provide more insights, we further investigate the proposed bi-directional adaptation strategy on the LQ set of FF++. We use the DeepFake detector trained on the source domain without adaptation as the baseline. GRL [69] and MMD [44] are two classical domain adaptation methods. GRL uses the gradient reversal layer and MMD attempts to reduce the distance of the probability distributions between source and target domains. GRL is exactly the domain adaptation method adopted in FEAT [54]. We denote the forward adaptation method in our method as FA. In the top part of Table VII, we evaluate these three domain adaptation methods. It can be seen that adding GRL, MMD or FA can improve the performance in the target domain, and using our FA attains the highest performance gain.
TABLE VII: Effect of various adaptation settings.
@@
Then we take a further step to analyze the effect of backward adaptation in the bottom part of Table VII, where SD means that the self-distillation is used in backward adaptation, while Ent. represents that self-distillation is replaced with entropy minimization of samples, which is inspired by the method of [70]. We observe that only adding SD can notably improve the performance on the target domain, but its performance on the source domain is compromised, especially in NT $\rightarrow$ FS, with $77.53\%$ of +SD v.s. $87.61\%$ of Baseline. This is because self-distillation is effective on the target domain, but overlooks the adaptation learned in forward adaptation. It can also be seen that replacing SD with Ent. is a bad idea, as this strategy degrades the detection performance in both domains. By adding both our forward adaptation and backward adaptation components to the baseline, our method can reach a good balance of the detection performance in both source and target domains. This clearly demonstrates that the proposed bi-directional adaptation is highly effective.
@@ IV-E2 Various Amounts of Training Samples
This part investigates the data efficiency of the proposed method. We randomly select a proportion or percentage of the samples in target domain for training to simulate data-constrained scenarios. Fig. 7 depicts the performance of our method for the adaptation pairs of DF $\rightarrow$ F2F, DF $\rightarrow$ NT, and F2F $\rightarrow$ FS on the LQ set of FF++ using the percentage of the target samples ranging from 20% to 100%. As expected, increasing the percentage generally improves achievable performance. More importantly, noting the baseline+FA performance of 86.49 for DF $\rightarrow$ F2F from Table VII, it can be seen that our method can still considerably enhance the performance to over 90 even if only 20% of target samples are used. This indicates that our method can be utilized in the more restricted cases where the available target samples are insufficient.
TABLE VIII: Effect of frequency module.
@@ IV-E3 Effect of Frequency Module
The number of channels (C) and depth (D) of DCT blocks are two important factors for extracting frequency features. The C and D values used in our method are 768 and 4. We now investigate the achievable performance of our method using various C and D values on the LQ set of FF++. As can be seen from Table VIII, given depth $\text{D}=4$, increasing channels from $\text{C}=192$, 384 to 768 improves the achievable performance on both source and target domains. Likewise, with channels fixed to $\text{C}=768$, increasing depth from $\text{D}=2$, 3 to 4 improves the achievable performance on both domains. The results of Table VIII also validate that our choice of $\text{C}=768$ and $\text{D}=4$ is appropriate.
TABLE IX: Performance of using various feature extractors.
@@ IV-E4 Various Feature Extractors
As mentioned in the introduction section, the proposed framework is independent of the architecture. Table IX shows the performance of applying our framework on ResNet-50 [30], Xception [31], EfficientNet-B0 [32], EfficientNet-B4 [32], ViT-small [29] and ViT-base [29] on the LQ set of FF++. It can be seen that for each base network, our method significantly improves the performance on the target domain while maintaining a favorable performance on the source domain. This demonstrates that our proposed method is generically applicable.
@@ IV-E5 Added on State-of-the-art DeepFake Detection
[…]
TABLE X: Performance of adding our method on SBI.
Adaptation Setting SBI [18] SBI + Ours
(HQ) SBI $\rightarrow$ DF 85.88 99.00
(HQ) SBI $\rightarrow$ F2F 80.07 97.23
(HQ) SBI $\rightarrow$ NT 72.47 90.28
(HQ) SBI $\rightarrow$ FS 75.62 92.92
Avg. 78.51 94.86
(LQ) SBI $\rightarrow$ DF 80.70 88.93
(LQ) SBI $\rightarrow$ F2F 67.63 68.20
(LQ) SBI $\rightarrow$ NT 65.13 64.98
(LQ) SBI $\rightarrow$ FS 62.50 65.52
Avg. 68.99 71.91
IV-E6 Temperature $\tau$
We conduct experiments to investigate the influence of temperature in Eq. 4. As seen in Table XI, we observe that the model training with a large $\tau$ (e.g. $>$ 0.5) achieves stable performance. However, when $\tau$ is set less than 0.5, the result will get unstable.
TABLE XI: Effect of temperature $\tau$.
Sub-task 0.35 0.5 0.65 0.8
[…]
IV-E7 Teacher model updating strategy
[…]
TABLE XII: Effect of teacher model updating strategy.
Strategy (HQ) DF $\rightarrow$ F2F (HQ) FS $\rightarrow$ NT
EMA [77] 98.49 94.19
Baseline+BA (Ours) 99.08 96.27
IV-E8 Results on unseen dataset
We further conduct an experiment to evaluate the generalization performance on unseen dataset, where we make a comparison on two adaptation settings named FF++(HQ) $\rightarrow$ CelebDF and FF++(HQ) $\rightarrow$ DFDCP. As shown in Table XIII, We observe that our method can also achieve comparable generalization performance on Celeb-DF when the target domain is DFDCP.
TABLE XIII: Cross-dataset generalization comparison.
Adaptation Setting CelebDF
FF++(HQ) $\rightarrow$ CelebDF 90.22
FF++(HQ) $\rightarrow$ DFDCP 85.24
@@ V Conclusions
VI Acknowledgement
This work is supported by National Key Research and Development Program of China under grant number 2022ZD0117201. Yuezun Li is supported by China Postdoctoral Science Foundation under grant number 2021TQ0314 and grant number 2021M703036.